The retroreflecting optics has been mounted on the telescope all the two days due to the bad weather conditions.
1) AO system calibration and performance
The measurements of the 4 reconstructors at BIN1,2,3,4 have been repeated using a new modal basis having more averaged measurements from 4D. The calibration data required to go on sky has been completed even in this case. The performance of the 4 reconstructors in terms of SR has been found to be consistent with results given by numerical simulations. Moreover with this new base the overall system gain is found to be more robust and constant for the various reconstructors.
2) AO system SW
A test of the AO system operation using the AOS and IRC commands has been done. After some debugging of the code the commmand sequence to preset AO system, acquire refernce star and close the loop has been tested succesfully. The test has been done with the telescope in tracking ode at elevations differemt from 90 degrees. The total time required to issue and complete the three commands preset_ao, acquire_ref, start_ao using the irc commmands has been compute to be 2 minuts and 50 seconds. This test has been done using AO5 and not BP4. This is because the BP4 has been found not able to guide in simulation mode.
3) AO system open loop transfer function.
A first measurement of the system open loop transfer function has been done for a few low order modes. This in order to characterize the system perfomance more accurately. Data reduction is ongoing.
4) test of M2 astigmatism introduced with elevation.
The AO system has been set in closed loop and the telescope elevation angle has been changed from 90 to 26 degree. The astigmatism in the mirror commands has been measured as before. This new set of data includes more positions then the previous one to characterize hysteresis effects. See the attached plots for the data. Data reduction is ongoing.